OpenStudy (anonymous):

Geometric mean of 11 and 1331?

OpenStudy (anonymous):

\[\sqrt{11*1331}\] Geometric mean of 2 numbers is the square root of their product

OpenStudy (anonymous):

Our teacher wants us to put it in radical form. So would that mean that I would just put the product (121) into radical form?

OpenStudy (jamesj):

Just leave the answer as 121. Your teachers point was to write a precise expression for it. For example, the geometric mean of 2 and 4 is \[ \sqrt{8} = 2\sqrt{2} \] this is radical form, and doesn't involve any messy, imprecise decimal expansions.

OpenStudy (jamesj):

*teacher's

OpenStudy (anonymous):

121 can be reduced to 11 though. So should I leave my answer as 11?

OpenStudy (jamesj):

No, it can't here because the answer is NOT \[ \sqrt{121} \]

OpenStudy (anonymous):

Ohh, okay. Thank you.

OpenStudy (jamesj):

\[ \sqrt{11 \cdot 1331} = \sqrt{11 \cdot 11^3} = \sqrt{11^4} = 11^2 = 121 \]